Show that the Integrating factor of the equation (x^2+y^2+2x)dx + 2ydy=0 is e^x and its particular solution is x^2 + y^2=2e^1-x when x=y=1
1
Expert's answer
2020-09-09T18:52:26-0400


We have the differential equation:

"(x^2+y^2+2x)dx+2ydy=0" ...(1)

let f(x,y)="x^2+y^2+2x"

g(y)=2y

so

f'"_y(x,y)=2y"

g'(x)=0


using Integration factor method,

"\frac{du}{dx}=\frac{(f_y-g_x)u}{g}"

"\frac{du}{dx}=\frac{(2y-0)u}{2y}"

"\frac{du}{dx}=\frac{(2y)u}{2y}"

"\frac{du}{dx}=u"

on solving we get integrated factor

u=e"^x"

Multiply equation 1 by integrated factor

"e^x(x^2+y^2+2x)dx+2ye^xdy=0"

let M="(x^2+y^2+2x)e^x"

N="2ye^x"

"\frac{dM}{dy}=2ye^x"

"\frac{dN}{dx}=2ye^x"

Hence "\frac{dM}{dy}=\frac{dN}{dx}"

Solution of above equation is given by,

Let I(x,y) be a implict function

I(x,y)="\int M(x,y)dx"

="\int e^x(x^2+y^2+2x)dx"

="x^2e^x-2xe^x+2e^x+y^2e^x+2xe^x-2e^x"

="x^2e^x+y^2e^x" +f(y)


Let differentiate I(x,y) with respect to y

"\frac{dI}{dy}=" N(x,y)

"2ye^x+\frac{df}{dy}=2ye^x"

"\frac{df}{dy}=0
=constant"

I(x,y)="x^2e^x+y^2e^x+c"

at x=y=1

0=e+e+c

c+2e=0

c=-2e

so Particular solution becomes

"x^2e^x+y^2e^x-2e=0"

"x^2+y^2=2e^{1-x}" .

This is the required answer.